A box contains 24 transistors, 4 of which are defective. If 4 are sold at random, find the
following probabilities.
a. Exactly 2 are defective. (5 marks)
c. All are defective. (5 marks)
b. None is defective. (5 marks)
d. At least 1 is defective. (5 marks)
1
Expert's answer
2021-07-29T20:37:03-0400

a. Exactly 2 are defective:

"P(2) = C(2;4)\u00d7(1\/6)^2\u00d7(5\/6)^2 = 6\u00d7(1\/6)^2\u00d7(5\/6)^2 = 0.1157."

c. All are defective:

"P(4) = (1\/6)^4 = 0.00077."

b. None is defective:

"P(0) = (5\/6)^4 = 0.4823."

d. At least 1 is defective:

P(>=1) = 1 - P(0) = 1 - 0.4823 = 0.5177.